Write the point-slope form of the equation of the line described.
1)through: (3, -4), parallel to y=3x-4
2)through: (5,2), perpendicular to y=-5x+5
I posted 2 problems because thy're related and i didnt want to waste all 4 problems I could post, because it's a really big packet of review units.

Answer by scott8148(6628)

1) parallel means same slope ___ y - -4 = 3 (x - 3)

2) perpendicular means negative-reciprocal slope ___ y - 2 = 1/5 (x - 5)